The breath of a rectangle is three cm shorter than the length.If the perimeter pf the rectangle is 58 cm, find the length and breadth of the rectangle.
Answer by Cromlix(4381)

Hi there,
Make the length = x
Breadth = x - 3
Perimeter = 2 x Length + 2 x Breadth
58 = 2(x) + 2(x - 3)
58 = 2x + 2x - 6
Collect like terms
2x + 2x = 58 + 6
4x = 64
x = 16
Length = 16 cm
Breadth = 13 cm
Hope this helps :-)